Suppose Scarlett purchased a 3-year bond with the face
value of 20,000 in the primary market. The current risk-free interest rate was
0.25%, and a risk premium on that bond is 3%. A year later after collecting her
yearly coupon payment, Scarlett decided to sell that bond in the secondary market.
By that time, the economic situation has improved, and the risk-free interest rate
has risen to 1%. What is Scarlett’s rate of return for the one-year period she held
the bond?
1
Expert's answer
2020-10-16T07:31:40-0400

Rate of return of the bond = "\frac {End value of principal + coupon interest- Beginning value of principal}{ Beginning value of principal} \times 100"



Expected return after one year = 3+0.25= 3.25%


Coupon interest= "\frac {3.25}{100} \times 20000 = 650"


End value of principal = "1.0325 \times 20000= 20650"


Substitute the values into the above formula


Therefore "\frac {20650 + 650- 20000} {20000} \times 100"


Rate of return for the year = 6.5 %